All the Heat generated by a current of 2Ampere passing through a 6ohms resistor for 25seconds is used to evaporate 5g of liquid at its boiling point. What is the specific Latent Heat of the Liquid?
1
Expert's answer
2019-07-24T15:50:18-0400

The energy balance equation

"I^2Rt=Lm"

So, the specific latent yeat of the liquid

"L=\frac{I^2Rt}{m}"

"=\frac{(2\:\rm{A})^2\times 6\:\Omega\times 25\:\rm{s}}{0.005\:\rm{kg}}=120\:000\;\rm{J/kg}"